OpenStudy (anonymous):

*help requested* *fan and medal awarded* Solve: 4/5 y + 2 =6 5 7 9 11

OpenStudy (anonymous):

subtract 2 from both sides first

OpenStudy (anonymous):

and

OpenStudy (anonymous):

then divide both sides by 4/5

OpenStudy (anonymous):

okay. i got -1.2 and 4

OpenStudy (anonymous):

@CalebBeavers

OpenStudy (anonymous):

4/5 y=4 y=4 * 5/4 y=?

OpenStudy (anonymous):

\[\frac{4}{5}y = 4\] Multiply by 5 both sides: \[\frac{4}{\cancel{5}} \times \cancel{5}y = 4 \times 5 \implies \color{blue}{4y = 20}\]

OpenStudy (anonymous):

Now divide both the sides by 4 and get your y..